# General properties of the electric Penrose process

### v1

@@ Abstract
We consider the Penrose process with the charged particles in the Reissner-Nordström background. Let parent particle 0 decay to particles 1 and 2. With the assumption that all three particles move in the same plane, the exact formulas for characteristics of particles 1 and 2 in terms of those of particle 0 are derived. We concentrate on scenarios in which particle 1 and 2 are ejected along the trajectory of particle 0. It is shown that such scenarios correspond to the extrema of energies $E_{1}$ or $E_{2}$ of daughter particles with respect to the angular momentum $L_{1}$ or $L_{2}$ of one of them. We derive bounds on the values of angular momenta $L_{1}$ and $L_{2}$. We give classification of these scenarios and discuss their properties including decay in the near-horizon region. The results are reformulated in terms of velocities of daughter particles in the center of mass frame. The approach is applicable also to collisional Penrose process, in which a combination of particles 1 and 2 is considered as one effective particle. If the mass of particle 0 $m_{0}\rightarrow\infty$, the situation corresponds to the Bañados-Silk-West effect, the results agree with the ones known in literature before. In addition, we consider special cases when decay occurs in the turning point for one or all three particles.
@@ I Introduction
The Penrose process (hereafter PP) means that a parent particle 0 decays to two daughter ones 1 and 2 in such a way that one of them (say, 1) has a negative energy whereas particle 2 returns to infinity with an energy bigger than an initial energy $E_{0}$ of particle 0. This process becomes possible if in a space-time there exists the region (called ergoregion or ergosphere) inside which the component of the metric $g_{tt}>0$, if signature of a metric is chosen to be (-,+,+,+) pen , pen2 . Meanwhile, a counterpart of the original PP in the Reissner-Nordström (RN) background was found due to the properties of particle dynamics even in spite of the absence of such an ergoregion ruf , den . Instead, in this metric negative energies are possible in some region (called generalized ergoregion) whose border depends on the electric charge, angular momentum and mass of a particle.
@@
At present, we see a new wave of interest to the PP in new contexts. This includes the collisional version of it schrev , confined one conf , myconf , the relation between the so-called Bañados-Silk-West (BSW) effect ban and the PP, discussion of the possibility of getting unbounded  energies called super-Penrose (SPP) effect (see, e.g. is and references therein), etc. As a rule, the investigations of the BSW effect were based on the careful analysis of the vicinity of the horizon. This implies that the value of the lapse function $N\ll 1$. Meanwhile, there is another approach that is based on exact formulas describing decay. For neutral particles this was realized in centr . The meaningful difference between (i) collisions near a rotating black hole and (ii) a static charged one consists in that in case (i) there is an upper bound forbidding the SPP near black holes but there is no such a bound for (ii) rn , fh .
@@
The emphasis in ruf , den and the main part of subsequent works of this trend (see dad3 for review) was made on the properties of negative energy states for a given particle in the presence of the electromagnetic field for a concrete metric rufkerr . Meanwhile, we are going to focus on another aspect connected with the relation between an initial state and its products after decay. We elucidate how characteristics of particle 0 and those of particles 1 and 2 are related with each other, which scenarios are possible in principle and what are their properties. Although for a realistic process in astrophysics it is necessary to take into account both rotation and electromagnetic field that includes not only the electric but also the magnetic one dad1 , dad2 , luca , fifty (see also a useful list of references on the subject in win ), it looks reasonable at the first stage to consider a simplified problem and separate two different factors - rotation and electric charge. In the previous work eq we gave full classification of possible scenarios for the particle decay and PP with neutral particles in the rotating  background, provided all processes occur within the equatorial plane. In the present work, we consider a similar problem for electrically charged particles in the static black hole metric with the charge, i.e. the Reissner-Nodström (RN) one. It is worth noting that a similar problem was considered recently in ju but for weakly charged black holes only and one scenario whereas in our approach we take into account this charge exactly and give classification of all possible scenarios. What is more important, our results are qualitatively different (see discussion below).
@@
The paper is organized as follows. In Sec. II we list equations of particle motion in the RN background. In Sec. III we give the exact formulas that related characteristics of particle 0 and particles 1, 2. In Sec. IV we suggest classification of scenarios when particle 0 is ingoing. In Sec. V we make emphasis on a type of decay when particles 1 and 2 are ejected along the trajectory of particle 0. In Sec. VI we consider a special case than an escaping particle is massless. In Sec. VII we consider some properties of scenario I including bounds on the energy and angular momentum of an escaping particle and decay in the near-horizon region. In Sec. VIII we discuss the situation when the point of decay coincides with the turning point of radial motion for particle 0, particle 2 or with a common turning point of all three particles. In Sec. IX we derive the conditions necessary for the PP process in the scenarios under discussion. In Sec. X we reformulate our result kinematically, using the velocities of particles and their Lorentz gamma-factors. In Sec. XI we consider the scenario in which particle 0 is outgoing. In Sec. XII we display the relation between our approach and that used in literature for the description of the BSW effect before. In Sec. XIII we give the summary of main results.
@@ II Charged particles in the metric of Reissner-Nordström black hole
with $d\omega^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}.$ We will mainly deal with the RN metric, then $f\equiv N^{2}=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}=\frac{(r-r_{+})(r-r_{-})}{r^{2}}$. Here, $M$ is a black hole mass, $Q$ is its electric charge (for definiteness $Q>0$), $r_{+}=M+\sqrt{M^{2}-Q^{2}}$ being the event horizon radius, $r_{-}$ the inner horizon radius. Meanwhile, basic results are extendable to an arbitrary $f$ outside the event horizon.
@@ III Decay to two particles: general scenario
Let particle 0 move with decreasing $r$, so $\sigma_{0}=-1$. It is clear that combination $\sigma_{1}=\sigma_{2}=+1$ is impossible since this would contradict eq. (14). For definiteness, we assume that particle 1 moves after decay with $\sigma_{1}=-1$. Moreover, we imply that in the frame comoving with particle 0 (which is the center of mass frame of particles 1 and 2 in the point of decay), particle 1 moves in the inward direction as well. Meanwhile, $\sigma_{2}$ can have any sign. In the center of mass frame particle 2 moves in the outward direction with some velocity $v_{2}$. However, if we pass to the frame comiving with particle 0 that moves in the inward direction with the velocity $V_{2}$ in the static frame, both cases are possible: $\sigma_{2}=-1$ or $\sigma_{2}=+1$ depending on which value of the velocity is bigger (below we will discuss this issue in more detail).
@@
IV Classification of scenarios in general
@@ II $(\sigma_{2}=-1$, $h_{2}<0$, $\delta=-1)$
$X_{2}=\frac{1}{2\tilde{m}_{0}^{2}}\left(X_{0}\tilde{b}_{2}-P_{0}\delta\sqrt{\tilde{d}}\right)\text{.}$ (38)
III $(\sigma_{2}=-1$, $h_{1}\leq 0$, $\delta=+1)$
@@
The condition $\tilde{d}\geq 0$ requires that, according to (25), (26)
$L_{1,2}^{(-)}\leq L_{1,2}\leq L_{1,2}^{(+)}\text{,}$ (43)
where
V Ejection along trajectory: maximization/minimization of energy and classification of scenarios
@@
V.1 Scenario I, $\sigma_{2}=+1$, $\delta=-1$, $h_{2}=+1$
@@
$P_{2}=\frac{P_{0}}{2m_{0}^{2}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}}(X_{0}\sqrt{d}-b_{2}\sqrt{X_{0}^{2}\mathcal{-}N^{2}m_{0}^{2}})\text{,}$ (51)
@@ V.2 Scenario II
VI Massless case
A special case arises when particle 2 is massless since condition (67) is no longer valid if $m_{2}=0$. In turn, this means that scenario II fails. This can be explained as follows. In general, scenario II is realized when particle 2 is ejected in the outward direction in the frame comoving with particle 0 (that is also the center of mass frame for particles 1 and 2). However, when we pass to the stationary frame, it is drifted in the inward direction due to motion of particle 0 (see Sec. X below for discussion of local Lorentz transformations). However, if particle 2 is massless, it moves with a speed of light and inward motion of particle 0 cannot overcome outward motion of particle 2. As a result, particle 2 moves outwardly that corresponds to scenario I, not II.
@@ VI Massless case
From another hand, if we put $m_{1}=0$, this does not give rise to any difficulties since particle 1 moves inwardly in both frames, so scenario II is valid.
@@
VII Properties of scenario I
@@
VII.1 Upper bound on the angular momentum
@@
VII.2 Upper bound on the energy of escaping particle
@@
VII.3 Near-horizon limit
@@
The Penrose process with charged particles was also considered in ju where an additional assumption that the black hole charge $Q$ is small was made. Our results do not agree. The authors of ju , according to their eqs. (34), (35), obtained that the outgoing particle after decay near the horizon has a finite nonzero energy. Meanwhile, it follows from our formulas that in this case the energy of an escaping particle (if we neglect $Q$) tends to zero. If we take the charge $Q$ into account, in a similar way the difference $X_{2}=E_{2}-q_{2}\frac{Q}{r_{+}}$ tends to zero. Actually, the subtlety in the issue under discussion consists in the necessity to take into account correlation between dynamics and kinematics. This means that in scenario I (where particle 2 escapes) the sign before the radical in (49) comes with minus, not plus. Instead, one can consider particle 1 and finite $X_{1}$ with the sign plus but this particle falls in a black hole and does not escape.
One more attempt of considering the PP for charged particles was made in sh with a magnetic field $B$ taken into account. In our view, the results for efficiency $\eta=\frac{E_{2}}{E_{0}}$ m (in our notations) described by Eqs. (31) and (32) of Ref. sh are incorrect. In the neutral case $q=0$, $Q=0$ and $B=0$ they give $\eta=0$ instead of the Schwarzschild value $\eta_{Sch}$. Therefore, for the RN metric, it also does not reproduce the formulas like (106) of the present paper.
VIII Special type of scenario: decay in turning point
There are special cases to be considered separately. They arise when decay occurs in the turning point. (Hereafter, by turning point we imply for brevity a turning point for radial motion, a particle can have in general an angular moment and nonzero angular velocity.) It follows from the conservation law (14) that the point of decay cannot coincide with the turning point for two particles precisely. Either it is the turning point for (i) only one particle or (ii) for all three at once. It follows from (14) or (52), (65) that in scenario I, case (i) is possible in the case of the turning point for particle 2, not for particle 1. Now, we will discuss (i) and (ii) case by case.
VIII.1 Case (i), turning point for particle 2 (TP2)

### v2

@@ Abstract
We consider the Penrose process with the charged particles in the Reissner-Nordström (RN) background. Let parent particle 0 decay to particles 1 and 2. With the assumption that all three particles move in the equatorial plane, the exact formulas for characteristics of particles 1 and 2 in terms of those of particle 0 are derived. We concentrate on scenarios in which particle 1 and 2 are ejected along the trajectory of particle 0. It is shown that such scenarios correspond to the extrema of energies $E_{1}$ or $E_{2}$ of daughter particles with respect to the angular momentum $L_{1}$ or $L_{2}$. We derive bounds on the values of angular momenta $L_{1}$ and $L_{2}$. We give classification of these scenarios and discuss their properties including decay in the near-horizon region. We find that the maximum of efficiency is achieved on the horizon for some of these scenarios but not for all of them and with additional constraints on particle parameters. The results are reformulated in terms of velocities of daughter particles in the center of mass frame. The approach is applicable also to collisional Penrose process, in which a combination of particles 1 and 2 is considered as one effective particle. If the mass of particle 0 $m_{0}\rightarrow\infty$, the situation corresponds to the Bañados-Silk-West effect, the results agree with the ones known in literature before. In addition, we consider special cases when decay occurs in the turning point for one or all three particles. The formalism developed in this work has a model-independent character and applies not only to the RN metric.
@@ I Introduction
The Penrose process (hereafter PP) means that a parent particle 0 decays to two daughter ones 1 and 2 in such a way that one of them (say, 1) has a negative energy whereas particle 2 returns to infinity with an energy bigger than an initial energy $E_{0}$ of particle 0. This process becomes possible if in a space-time there exists the region (called ergoregion or ergosphere) inside which $g_{\mu\nu}\xi^{\mu}\xi^{\nu}>0$, where $g_{\mu\nu}$ are metric coefficients and $\xi$ is the Killing vector responsible for time translation (the signature of a metric is chosen to be (-,+,+,+)) pen , pen2 . Meanwhile, a counterpart of the original PP in the Reissner-Nordström (RN) background was found due to the properties of particle dynamics even in spite of the absence of such an ergoregion ruf , den . Instead, in this metric negative energies are possible in some region (called generalized ergoregion) whose border depends on the electric charge, angular momentum and mass of a particle.
More recently, we have seen a new wave of interest to the PP in new contexts. This includes the collisional version of it schrev , confined one conf , myconf , the relation between the so-called Bañados-Silk-West (BSW) effect ban and the PP, the PP for spinning particles ext , binaries bin , pp2bh , the PP in the background of the Vaidya space-time ppv . In case of getting formally unbounded  energies, the PP is called super-Penrose (SPP) effect (see, e.g. is and references therein). The subtleties connected with the difference between power and efficiency of the PP was discussed in trade .
@@
As a rule, the investigations of the BSW effect were based on the careful analysis of the vicinity of the horizon. If we characterize the proximity to the horizon by the value of the lapse function $N$, this implies that there is a small parameter $N\ll 1$. Meanwhile, there is another approach that is based on exact formulas describing decay. For neutral particles this was realized in centr . The meaningful difference between (i) collisions near a rotating black hole and (ii) a static charged one consists in that in case (i) there is an upper bound forbidding the SPP near black holes but there is no such a bound for (ii) rn , fh .
The emphasis in ruf , den and the main part of subsequent works of this trend (see dad3 for review) was made on the properties of negative energy states for a given particle in the presence of the electromagnetic field for a concrete metric rufkerr . Meanwhile, we are going to focus on another aspect connected with the relation between an initial state and its products after decay. We elucidate how characteristics of particle 0 and those of particles 1 and 2 are related with each other, which scenarios are possible in principle and what are their properties.
@@
Although for a realistic process in astrophysics it is necessary to take into account both rotation and electromagnetic field that includes not only the electric but also the magnetic one dad1 , dad2 , luca , fifty , uniform (see also a useful list of references on the subject in win ), it looks reasonable at the first stage to consider a simplified problem and separate two different factors - rotation and electric charge. In the previous work eq we gave full classification of possible scenarios for the particle decay and PP with neutral particles in the rotating  background, provided all processes occur within the equatorial plane. In the present work, we consider a similar problem for electrically charged particles in the static black hole metric with the charge, i.e. the Reissner-Nodström (RN) one. It is worth noting that a similar problem was considered recently in ju but for weakly charged black holes only and one scenario whereas in our approach we take into account this charge exactly and give classification of all possible scenarios. What is more important, our results are qualitatively different (see discussion below).
Although the main area of application is the RN metric, practically all formulas do not use its particular form. This means that our results apply to the wide class of spherically symmetric static metrics described by eq. (1) below. Although they do not apply to axially symmetric stationary ones directly, we hope that general idea is useful for them as well, even under the presence of the electromagnetic field. The reason is that our scheme is quite general. (i) We do not restrict ourselves by a particular scenario in which decay occurs in a turning point of radial motion but build classification scheme of all possible scenarios, (ii) we find explicitly the expression for a velocity of fragments after decay, (iii) we trace the relation between the electromagnetic version dad0 of the Wald approach wald and particle dynamics of decay, thus relating the velocities and masses, (iv) we find explicitly the angular momenta of daughter particles. We will see that the roles of rotation and electric charge are very different and some scenarios of the PP forbidden inside the ergosphere of the rotating metric eq are allowed in the generalized ergosphere of the static charged metric. Thus the present work and the previous one eq constitute the mutually complementary pair that we hope to combine in a unified picture later.
A separate question is the dependence of the efficiency of the process on the position of the decay point. There is a popular aproach in which this pint is chosen as close to the horizon as possible, with point of decay coinciding with a turning point of particle 0 (see, e.g. Sec. 5.1 of dad1 ). Meanwhile, we show that this is not always so and depends strongly on a type of scenario the list of which we discuss carefully case by case. Bearing in mind future application to astrophysical process, this seems to be important enough since it alows to describe different cases without additional arificial assumptions about a particular type of decay.
The paper is organized as follows. In Sec. II we list equations of particle motion in the RN background. In Sec. III we give the exact formulas that related characteristics of particle 0 and particles 1, 2. In Sec. IV we suggest classification of scenarios when particle 0 is ingoing. In Sec. VI we make emphasis on a type of decay when particles 1 and 2 are ejected along the trajectory of particle 0. In Sec. VII we consider a special case than an escaping particle is massless. In Sec. VIII we consider some properties of scenario I including bounds on the energy and angular momentum of an escaping particle and decay in the near-horizon region. In Sec. IX we discuss the situation when the point of decay coincides with the turning point of radial motion for particle 0, particle 2 or with a common turning point of all three particles. In Sec. X we derive the conditions necessary for the PP process in the scenarios under discussion. In Sec. XI we reformulate our result kinematically, using the velocities of particles and their Lorentz gamma-factors. In Sec. XII we consider the scenario in which particle 0 is outgoing. In Sec. XIII we consider the conditions under which the maximum of the efficiency is achieved on the horizon depending on scenario. In Sec. XIV we summarize the main features of the considered scenarios with short comments. In Sec. XV we compare our results with some other ones recently published and explain the origin of discrepancy. In Sec. XVI we display the relation between our approach and that used in literature for the description of the BSW effect before. In Sec. XVII we give the summary of main results. In Appendix we derive some useful inequalities relevant for the PP near the horizon.
@@ II Charged particles in the metric of Reissner-Nordström black hole
with $d\omega^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}.$ We will mainly deal with the RN metric, then $f\equiv N^{2}=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}=\frac{(r-r_{+})(r-r_{-})}{r^{2}}$. Here, $M$ is a black hole mass, $Q$ is its electric charge (for definiteness $Q>0$), $r_{+}=M+\sqrt{M^{2}-Q^{2}}$ being the event horizon radius, $r_{-}$ $=M-\sqrt{M^{2}-Q^{2}}$ the inner horizon radius. Meanwhile, basic results are extendable to an arbitrary $f$ outside the event horizon.
@@ III Decay to two particles: general scenario
Let particle 0 move with decreasing $r$, so $\sigma_{0}=-1$. It is clear that combination $\sigma_{1}=\sigma_{2}=+1$ is impossible since this would contradict eq. (14). For definiteness, we assume that particle 1 moves after decay with $\sigma_{1}=-1$. Moreover, we imply that in the frame comoving with particle 0 (which is the center of mass frame of particles 1 and 2 in the point of decay), particle 1 moves in the inward direction as well. Meanwhile, $\sigma_{2}$ can have any sign. In the center of mass frame particle 2 moves in the outward direction with some velocity $v_{2}$. However, if we pass to the frame comoving with particle 0 that moves in the inward direction with the velocity $V_{2}$ in the static frame, both cases are possible: $\sigma_{2}=-1$ or $\sigma_{2}=+1$ depending on which value of the velocity is bigger (below we will discuss this issue in more detail).
@@ III.1 Unconditional bounds on mass
The same condition requires that, according to (25), (26)
$L_{1,2}^{(-)}\leq L_{1,2}\leq L_{1,2}^{(+)}\text{,}$ (28)
where
If one takes into account the definition (7) and the conservation law (11), one can obtain by straightforward algebraic manipulations that
$m_{0}\geq m_{1}+m_{2}$ (30)
as well.
One can also use the conservation law for four-momenta in the covariant form
$p_{0}^{\mu}=p_{1}^{\mu}+p_{2}^{\mu}\text{.}$ (31)
Then, taking the square and taking into account that for future-directed four-vectors $p_{1\mu}p_{2}^{\mu}<0$, it is easy to obtain
$m_{0}^{2}\geq m_{1}^{2}+m_{2}^{2}\text{.}$ (32)
This is just eq. (4.11) in dad1 and eq. (3.11) in dad2 , written in our notations. Meanwhile, eq. (30) is more tight than (32). It can be derived if one uses the property $p_{1\mu}p_{2}^{\mu}<-m_{1}m_{2}$ which is more tight than $p_{1\mu}p_{2}^{\mu}<0$ used in dad1 , dad2 .
IV Classification of scenarios with respect to parameters in general
@@ II $(\sigma_{2}=-1$, $h_{2}=-1$, $\delta=-1)$
$X_{2}=\frac{1}{2\tilde{m}_{0}^{2}}\left(X_{0}\tilde{b}_{2}-P_{0}\sqrt{\tilde{d}}\right)\text{.}$ (43)
III $(\sigma_{2}=-1$, $h_{1}=0$ or $h=-1$, $\delta=+1)$
@@
V Types of scenarios and their meaning
Formulas of the previous Section describe general relations between a parent particle and daughter ones. Meanwhile, for physical purposes, we are interested in more concrete scenarios. The most simple and popular one consists in such a decay of particle 0 to 1 and 2 that all three particles are located in the corresponding turning points. Meanwhile, in any real process, one cannot guarantee the validity of this condition and, therefore, one is led to considering more expanded set of possible scenarios. In the first place, this concerns the situation when all three particles are moving and the products of decay are ejected along the trajectory. This is of special interest since this leads to the most effective (for a given  state if particle 0) kind of process. (This will be proven in the next Section.) In turn, here there are different cases depending on how a daughter particle move after decay, whether it moves towards a black hole or towards infinity. There are also intermediate cases when some of particles (but not all of them) are in their turning points.
The key ingredient of the Penrose process is the existence of the ergoregion (in the standard PP) or generalized ergoregion (in the electric version of PP), where particle energy can be, in principle, negative. In this respect it was noticed earlier eq that for a rotating black hole some kinds of decay, which seem to be compatible with the existence of negative states for an individual particle, were forbidden in the ergoregion. It is of interest to elucidate, whether or not this happens for charged particles in the generalized ergoregion of a static charged black hole.
More precisely, we consider the following scenarios.
I: particle 2 moves in the outward direction, particle 1 moves in the inward one. Further, particle 1 can fall in a black hole while particle 2 can escape.
II: particles 1 and 2 move in the same direction as particle 0, i.e. inward direction.
In scenarios I and II it is implied that before decay particle 0 moved in the inward direction, so it follows from the momentum conservation that both particles 1 and 2 cannot move after decay in the outward direction.
B (”bounce”). Particle 0 bounces back from the potential barrier, moves in the outward direction and decays afterwards. Preliminarily, this scenario gives us the most efficient possibility when a daughter particle moves in the same direction as particle 0. Formally, this is similar to scenario II but in scenario II both particles move towards a black hole whereas in scenario B they start to move to infinity.
For completeness, we take into account also the existence of a turning point.
TP2: point of decay of particle 0 coincides with the turning point of decay of particle 2.
TP0: point of decay of particle 0 coincides with its turning point but not with the turning points of particles 1 and 2.
TP3: decay occurs in the common point of decay for all three particles.
Below, we consider all these scenarios case by case. Bearing in mind potential applications to processes near a black hole, of primary interest is the near-horizon behavior of particles in these processes. The results are collected in Sec. XIV below where we also indicate near-horizon properties. A special questin about the role of a horizon in achieving the maximum of efficiency is considered in Sec. XIII.
VI Ejection along trajectory: maximization/minimization of energy and classification of scenarios
@@
VI.1 Scenario I, $\sigma_{2}=+1$, $\delta=-1$, $h_{2}=+1$
@@
$P_{2}=\frac{P_{0}}{2m_{0}^{2}\sqrt{X_{0}^{2}-m_{0}^{2}N^{2}}}(X_{0}\sqrt{d}-b_{2}\sqrt{X_{0}^{2}\mathcal{-}m_{0}^{2}N^{2}})\text{,}$ (54)
@@ VI.2 Scenario II
VII Massless case
A special case arises when particle 2 is massless since condition (70) is no longer valid if $m_{2}=0$. In turn, this means that scenario II fails. This can be explained as follows. In general, scenario II is realized when particle 2 is ejected in the outward direction in the frame comoving with particle 0 (that is also the center of mass frame for particles 1 and 2). However, when we pass to the stationary frame, it is drifted in the inward direction due to motion of particle 0 (see Sec. XI below for discussion of local Lorentz transformations). However, if particle 2 is massless, it moves with a speed of light and inward motion of particle 0 cannot overcome outward motion of particle 2. As a result, particle 2 moves outwardly that corresponds to scenario I, not II.
@@ VII Massless case
From another hand, if we put $m_{1}=0$, this does not give rise to any difficulties since particle 1 moves inwardly in both frames, so scenario II is valid as well.
VIII Properties of scenario I
@@
VIII.1 Upper bound on the angular momentum
@@
VIII.2 Upper bound on the energy of escaping particle
@@
VIII.3 Near-horizon limit
@@
IX Special type of scenario: decay in turning point
There are special cases to be considered separately. They arise when decay occurs in the turning point. (Hereafter, by turning point we imply for brevity a turning point for radial motion, a particle can have in general an angular moment and nonzero angular velocity.) It follows from the conservation law (14) that the point of decay cannot coincide with the turning point for two particles precisely. Either it is the turning point for (i) only one particle or (ii) for all three at once. It follows from (14) or (55), (68) that in scenario I, case (i) is possible if there exists a turning point for particle 2, not for particle 1. Now, we will discuss (i) and (ii) case by case.
IX.1 Case (i), turning point for particle 2 (TP2)